The achicago LaTeX package

Chicago Manual A-style (but author-date) citations

          Matt Swift <swift@alum.mit.edu>

Version: 1 Date: 1995/11/27

Documentation revision: 1999/03/01

Abstract

A bibliography style based on The Chicago Manual of Style. The style is mostly style A, but borrows the author-date reference system from style B. Section 16.25 of the manual permits this combination. Requires accompanying BibTEX bibliography style, achicago.

Contents

I Discussion 2

1 General 2 2 Notes on the future of this package and bibstyle 2 3 Usage 3 4 Some technical notes 4 5 Credits 4

II Implementation 5

6 Version control 5

Part I

Discussion

1 General

Warning: This documentation is spotty because I want to release this in time to get on TEXLive 4. Sorry! Despite appearances in this file, this package really does work. My testbed has been a long document with about 350 citations and a 12-page bibliography.

Warning: Annotation was working once, but I haven’t toyed with it in a long time.

This package is designed for use with the achicago BibTEX bibliography style. For the moment, you pretty much have to commit to using this bibstyle-package combination when you write your sources. This needn’t be true, and one day I am going to create a series of mappings from other common bibstyles that are conceptually similar, such as the harvard styles. The user commands are slightly different, but it should be the case that either set of user commands can be the front end for either bibstyle.

Here in this documentation you will read about the commands you will use in your L^{A}TEX source file to make citations, and what the citations look like. Documentation of the achicago bibstyle itself is in the file achicago-bst.dvi. There you can read about what the References section, the actual book list, looks like. There are also some new fields recognized and other information you might want to know that relates to your BibTEX bibliography database file (bib file). You may also wish to look at the titles package (also in the Frankenstein bundle), which can be very helpful in typesetting various styles of titles properly, even when nested. The achicago package already requires the titles package, so those commands are always there if you want to use them.

2 Notes on the future of this package and bibstyle

In the future I hope to document more closely the Manual’s principles wrt each entry type, the many choices given by The Chicago Manual.

Untested and indeed hardly testable nature of the subject, with all its many permutations and difficult special bibliographic cases. Feedback is very welcome, especially with citations from the Manual.

FUTURE One thing I realize now is that I’ve kept the same user interface I inherited, which I don’t think corresponds with any other popular style. Besides refinement, this is the next major step in the future of this bibstyle-package.

3 Usage

Some of this dox is spotty, some redundant. Here are the basic inline citations that you can conjure with this package:

%   Citation format:
%             (author-last-name year)
%             (author-last-name and author-last-name year)
%             (author-last-name et al. year)
%             (author-last-name)
%             author-last-name
%             author-last-name (year)
%             (author-last-name and author-last-name)
%             (author-last-name et al.)
%             (year) or (year, year)
%             year or year, year
%

Here are the various citing commands, and what they look like. Presently the variable which controls the number of names before switching to “et al.” in the abbreviated author list is 2, but this can be made 3 or more easily with a simple change to a constant in ‘format.short.label.names’. This unfortunately cannot be controlled by TEX. I don’t think spacing after the period in “et al.” is handled correctly: FIX. Also, what about the situation when you end a sentence with something ending with “et al.”—in this case you don’t want to reproduce the period.

\cite{key}

full author list and year (Brown 1978; Jarke, Turner, and Stohl 1985)

\citeNP{key}

full author list and year, no parentheses Brown 1978; Jarke, Turner, and Stohl 1985

\citeA{key}

full author list (Brown; Jarke, Turner, and Stohl)

\citeANP{key}

full author list, without parentheses Brown; Jarke, Turner, and Stohl

\citeN{key}^{1}

full author list and year with only the year in parentheses, i.e., in the form of a noun Brown (1978) states that …

\shortcite{key}

abbreviated author list and year (Brown 1978; Jarke et al. 1985)

\shortciteNP{key}

abbreviated author list and year, no parentheses Brown 1978; Jarke et al. 1985

\shortciteA{key}  abbreviated author list  (Brown; Jarke et al.)

\shortciteANP{key}

abbreviated author list, no parentheses

Brown; Jarke et. al

\shortciteN{key}^{1}

abbreviated author list and year, in form of a noun Jarke et al. (1985) state that …

\citeyear{key}  year in parentheses  (1978, 1985)
\citeyearNP{key}  year no parentheses  1978, 1985

To do: Should I use a warning in case more than one key is given to a command that shouldn’t have them?

4 Some technical notes

To do: See §16.14 for issues to do with multiple citations.

This BibTEX style has support for abbreviated author lists and for year-only citations. This is done by having the citations actually look like:

\citeauthoryear{full-author-info}{abbrev-author-info}{year}

5 Credits

I wrote this package incorporating chicago\{,a\}.bst and filechicago.sty (v4, 92/8). By now I’ve almost totally rewritten it. It was a great mess because it had been based on newapa.bst which had been based on several other bibstyles, and each author had really just hacked up what was there in the previous bibstyle apparently without fully understanding it. Now, of course, it is a shining example of clarity and efficiency.

1This command should only be used for one key.

Part II

Implementation

6 Version control

\fileinfo \DoXUsepackagE \HaveECitationS \fileversion \filedate \docdate \PPOptArg These definitions must be the first ones in the file.
 1 \def\fileinfo{Chicago Manual A-style (but author-date) citations}
 2 \def\DoXPackageS {achicago}
 3 \def\initelyHavECitationS {}
 4 \def\fileversion{v1}
 5 \def\filedate{1995/11/27}
 6 \def\docdate{1999/03/01}
 7 \edef\PPOptArg {%
 8 \filedate\space \fileversion\space \fileinfo
 9 }

If we’re loading this file from a \ProcessDTXFile command (see the compsci package), then \JusTLoaDInformatioN will be defined; othewise we assume it is not (that’s why the FunkY NamE).

If we’re loading from \ProcessDTXFile, we want to load the packages listed in \DoXPackageS (needed to typeset the documentation for this file) and then bail out. Otherwise, we’re using this file in a normal way as a package, so do nothing. \DoXPackageS, if there are any, are declared in the dtx file, and, if you’re reading the typeset documentation of this package, would appear just above. (It’s OK to call \usepackage with an empty argument or \relax, by the way.)

10 \makeatletter% A special comment to help create bst files.  Don't change!
11 \@ifundefined{JusTLoaDInformatioN} {%
12 }{% ELSE (we know the compsci package is already loaded, too)
13 \UndefineCS\JusTLoaDInformatioN
14 \SaveDoXVarS
15 \eExpand\csname DoXPackageS\endcsname\In {%use \csname in case it's undefined
16  \usepackage{#1}%
17 }%
18 \RestoreDoXVarS
19 \makeatother
20 \endinput
21 }% A special comment to help create bst files.  Don't change!

Now we check for L^{A}TEX2e and declare the LaTeX package.

22 \NeedsTeXFormat{LaTeX2e}
23 \ProvidesPackage{achicago}[\PPOptArg]
24 \RequirePackage{moredefs,titles,blkcntrl}
25 \newboolean{Annotate}
26 \newcommand\annotate {%
27 \Annotatetrue

28 }

29 \newcommand\noannotate {%
30 \Annotatefalse

31 }

\citework is supposed to be a general command for citing things declared with \newwork in the abbrevs package. It has one optional and one required argument so that it is parallel with the other citing commands, but I cannot see any use for it without the optional argument. Environments can exert complete control over how this macro looks by resetting the three parameters. The default will look good outside all environments, in running text.

The second argument is expected to be something defined with \newwork. Needs modification to handle the empty optional arg. Watch interfering with

things surrounding macros might have set. \relax’s are intentionally left out to let constructions like \csname … \endcsname [eh? FIX] work on the arguments.

\PreCiteWork
\PostCiteWork
32 \providesavebox\sc@box@a
33 \newcommand\PreCiteWork {%
34 (\csname%

35 }

36 \newcommand\PostCiteWork {%
37 \end{lrbox}\usebox{\sc@box@a})%

38 }

We don’t want to be unbreakable here, but we want a high penalty. We absolutely do not want to break the number range, so we put it in an lrbox.

I think comma is better, even though it might seem fussy, because it is better parallel with the way \cite works with an optional page argument: the convention is that page numbers come after commas.

39 \newcommand\MidCiteWork {%
40 \endcsname,\penalty9000\ \begin{lrbox}{\sc@box@a}%

41 }

42 \newcommand\citework [2] {%
43 \PreCiteWork #2\MidCiteWork #1\PostCiteWork

44 }

% The {} fools abbrevs.dtx into not adding an extra space
% \newcommand\MidCiteWork {%
%   \endcsname{}\penalty9000\ \begin{lrbox}{\sc@box@a}%
% }
%

To do: Make citework* with no parentheses, or other alternative.

FIX: When the ? is placed there, there are two left parens, one right.

We want the remaining macros in this section to be available in their own piece.

To do: is \PreChunk the only dependence on blkcntrl? Should make this not necessary if so.

45 \newcommand\PreAnnotation {%
46 \PreChunk

47 }

This will make the definition of the thebibliography environment in classes. dtx do the right thing. FIX: not defined in letter class?

48 \defcommand\@openbib@code {%
49 \advance\leftmargin\bibindent
50 \itemindent -\bibindent
51 \listparindent \itemindent
52 \parsep \z@

53 }

54 \let\newblock\relax

This doesn’t work at the beginning, for some reason. The auxfiles are not set up right?

55 \AtEndDocument {%
56 \bibliographystyle{achicago}%

57 }

The achicago bibliography style will insert some macros that are not defined by L^{A}TEX, and some that must have new meanings. They are: \citeN, \SCcite,

\SCduplicate, \begin{SCannotation}, \end{SCannotation}.

Some of these commands should properly have @ in their names, but @- commands cannot appear in the bbl file. As a compromise, the names have the prefix SC.

\SCduplicate The argument will contain the actual ‘long.label’ that is a duplicate, in case it might ever be of use. But for now, we just want to replace duplicates with 3-em dashes.

FIX: OK, we’re trying to do it so that we print the argument if it’s the first \SCduplicate on the page, using the afterpage package. Problem is, we process sometimes one or two more \bibitem’s than make it onto the page. This is going to be next to impossible, even if you change the \output routine. So comment it out, including requiring afterpage.

Instead, we define \topitem, which can be inserted manually into the bbl file right before the \bibitem that occurs with an \SCduplicate at the top of a final page. The \topitem command causes the immediately following (and only that one) \SCduplicate to print its argument.

58 %\RequirePackage{afterpage}
59 %\newboolean{@sc@duplicatedash@}
60 %\@sc@duplicatedash@true
61 %\SaveCS\bibitem
62 %\newcommand\sc@bibitemfirst {%
63 %  \typeout{in bibitemfirst  [\the\foo]}%
64 %  \afterpage {%
65 %    \typeout{in afterpage [\the\foo]}%
66 %    \global\@sc@duplicatedash@false
67 %    \global\let\bibitem\sc@bibitemfirst

68 %  }%
69 %  \let\bibitem\MDSavedbibitem
70 %  \bibitem
71 %}
72 %\let\bibitem\sc@bibitemfirst

73

74 \newcommand*\sc@duplicate@dash [1] {%
75 %  \if@sc@duplicatedash@
76  \sc@mmmdash
77 %  \else
78 %    #1%
79 %  \fi
80 %  \@sc@duplicatedash@true

81 }

82

83 \newcommand\sc@duplicate@arg [1] {%
84 #1%
85 \let\SCduplicate\sc@duplicate@dash

86 }

87 \newlet\SCduplicate\sc@duplicate@dash
88 \newcommand\topitem {%
89 \let\SCduplicate\sc@duplicate@arg

90 }

\sc@mmmdash A 3em-dash.
91 \newcommand\sc@mmmdash {%
92 \rule[.6ex]{3em}{.03ex}%

93 }

\PreAnnotation \sc@begingobble \sc@endgobble SCannotation This sets up the SCannotation environment. When the boolean \IfAnnotate is
false, we gobble everything between \begin{SCannotation} and \end{SCannotation}.

To do: These should be comment and endcomment, but we have to be careful how we define them. An lrbox required balanced text inside.

94 \newcommand\sc@begingobble {%
95 \relax\begin{lrbox}{\sc@box@a}%

96 }

97 \newcommand\sc@endgobble {%
98 \relax\end{lrbox}\sbox\sc@box@a{}%

99 }

To do: I shouldn’t define annote in terms of quotation, we should copy a standard one here.

100 \newenvironment{SCannotation} {%
101  \ifAnnotate
102  \let\PreQuotation\PreAnnotation
103  \relax\begin{quotation}%
104  \else
105  \sc@begingobble
106  \fi

107 }{%
108 \ifAnnotate
109  \relax\end{quotation}%
110 \else
111  \sc@endgobble
112 \fi

113 }

\SCcite \sc@one@three \sc@two@three \SCcite is what achicago produces. Its args are ‘long.label’, ‘short.label’, and ‘year.label’.

114 \ReserveCS\SCcite

115

116 \newcommand\sc@one@three [3] {#1}
117 \newcommand\sc@two@three [3] {#2}

This spits out arguments #1 and #3 with #2 intervening iff \@tempswatrue.

  FIX -- what's going on
% \newcommand\sc@citeformat [3]
%  {\def\@cite#1#2{%
%    #1\if@tempswa#2\fi#3}
%

\PreCite \PostCite Should have one single command be what each of the following executes. Its arguments determine what all the other stupid things do. hunh? FIX

118 \newcommand\PreCite {%

119 (% 120 }

121 \newcommand\PostCite {%

122 )% 123 }

\cite \citeNP \citeA \citeN \citeANP The way this works is: \\@cite is called once for each citing command, and \SCcite is called once for each key. The results are spaced by either semicolons (\\sc@cite@sc) or commas (\\sc@cite@comma), and these become argument #1

for \\@cite.
   \PreCite1[, 2]\PostCite  \sc@citeformat{\PreCite##1}{, ##2}{\PostCite}
124 \def\cite {%
125 \defcommand*\@cite [2] {%
126  \PreCite ##1\if@tempswa , ##2\fi\PostCite

127 }%

128 \defcommand*\SCcite [3] {%
129  ##1\ ##3%

130 }%

131 \sc@cite@sc

132 }

133 % 1[, 2]    \sc@citeformat{##1}{, ##2}{}
134 \newcommand*\citeNP {%
135 \defcommand*\@cite [2] {%

136  ##1\if@tempswa , ##2\fi

137 }%

138 \defcommand*\SCcite [3] {%
139  ##1\ ##3%

140 }%

141 \sc@cite@sc

142 }

143 % 1[, 2])  \sc@citeformat{##1}{, ##2}{)}
144 % \newcommand*\citeN {%
145 %   \defcommand*\@cite [2] {%
146 %     ##1\if@tempswa , ##2\fi
147 %   }%
148 %   \defcommand*\SCcite [3] {%
149 %     ##1\ \PreCite##3\PostCite
150 %   }%
151 %   \sc@cite@comma

152 % }

‘long.label’ (‘year’, optarg)

153 \newcommand*\citeN {%
154 \defcommand*\@cite [2] {%
155  ##1\if@tempswa , ##2\fi\PostCite

156 }%

157 \defcommand*\SCcite [3] {%
158  ##1\ \PreCite ##3%

159 }%

160 \sc@cite@comma

161 }

162 \newlet\UnexpandableProtect\@unexpandable@protect
163 % \PreCite1[, 2]\PostCite  \sc@citeformat{\PreCite ##1}{, ##2}{\PostCite}
164 \newcommand*\citeA {%
165 \defcommand*\@cite [2] {%
166  \PreCite ##1\if@tempswa , ##2\fi\PostCite

167 }%

168 \let\SCcite\sc@one@three
169 \sc@cite@sc

170 }

171 % 1[, 2]    \sc@citeformat{##1}{, ##2}{}
172 \newcommand*\citeANP {%
173 \defcommand*\@cite [2] {%
174  ##1\if@tempswa , ##2\fi

175 }%

176 \let\SCcite\sc@one@three\sc@cite@sc

177 }

\shortcite
\shortciteNP
\shortciteN
\shortciteA
\shortciteANP

178 \newcommand*\shortcite {%
179 \defcommand*\@cite [2] {%
180  \PreCite ##1\if@tempswa , ##2\fi\PostCite}%
181 \defcommand*\SCcite [3] {%

182  ##2\ ##3%

183 }%

184 \sc@cite@sc

185 }

186 \newcommand*\shortciteNP {%
187 \defcommand*\@cite [2] {%
188  ##1\if@tempswa , ##2\fi

189 }%

190 \defcommand*\SCcite [3] {%
191  ##2\ ##3%

192 }%

193 \sc@cite@sc

194 }

195 \newcommand*\shortciteN {%
196 \defcommand*\@cite [2] {%
197  ##1\if@tempswa , ##2\PostCite\else\PostCite\fi}%
198 \defcommand*\SCcite [3] {%
199  ##2\ \PreCite ##3%

200 }%

201 \sc@cite@comma

202 }

203 \newcommand*\shortciteA {%
204 \defcommand*\@cite [2] {%
205  \PreCite ##1\if@tempswa , ##2\fi\PostCite

206 }%

207 \let\SCcite\sc@two@three
208 \sc@cite@sc

209 }

210 \newcommand*\shortciteANP {%
211 \defcommand*\@cite [2] {%
212  ##1\if@tempswa , ##2\fi

213 }%

214 \let\SCcite\sc@two@three\sc@cite@sc

215 }

\citeyear \citeyearNP
216 % \PreCite1[, 2]\PostCite  \sc@citeformat{\PreCite ##1}{, ##2}{\PostCite}
217 \newcommand*\citeyear {%
218 \defcommand*\@cite [2] {%
219  \PreCite ##1\if@tempswa , ##2\fi\PostCite

220 }%

221 \defcommand*\SCcite [3] {%
222  ##3%

223 }%

224 \sc@cite@comma

225 }

226 % 1[, 2]    \sc@citeformat{##1}{, ##2}{}
227 \newcommand*\citeyearNP {%
228 \defcommand*\@cite [2] {%
229  ##1\if@tempswa , ##2\fi

230 }%

231 \defcommand*\SCcite [3] {%
232  ##3%

233 }%

234 \sc@cite@comma

235 }

\sc@citesep
\sc@cite@sc
\sc@cite@comma

236 \ReserveCS\sc@citesep
237 \newcommand\sc@cite@sc {%
238 \let\sc@citesep ;%
239 \sc@cite

240 }

241 \newcommand*\sc@cite@comma {%
242 \let\sc@citesep ,%
243 \sc@cite

244 }

\sc@cite This command executes \\b@foo for every \foo in the list of cited labels, and

separates them by arg #1.

There has got to be a more elegant solution to this whole thing. FIX

245 \newcommand*\sc@cite {%
246 \@ifnextchar [ {%
247  \@tempswatrue
248  \sc@@cite
249  }{% ELSE
250  \@tempswafalse
251  \sc@@cite[]%

252 }% 253 }

To do: handle reserving names

254 \providecommand\@writeaux {%
255 \immediate\write\@auxout

256 }

257 \NewName*{sc@@cite} {[#1]#2} {% args: [optarg] label % optarg MANDATORY
258 \if@filesw
259  \@writeaux{\string\citation{#2}}%
260 \fi
261 \@cite{%
262  \InitCS\sc@t@a
263  \@for\sc@label:=#2\do {%
264  \sc@t@a
265 %      \let\sc@t@a\sc@citesep
266  \def\sc@t@a {\sc@citesep\ }% add space
267  \@ifundefined{b@\sc@label} {%
268  {\bfseries ?}%
269  \@warning{Citation `\sc@label' on page \thepage\space undefined}%
270  }{% ELSE
271  \@nameuse{b@\sc@label}%

272 }% 273 }%

274 }{#1}%

275 }

\bibindent Indent second and subsequent lines of bibliographic entries.

276 \setlength\bibindent{1.5em}

thebibliography There is no openbib option. The definitions of \newblock and \@ biblabel are kept local in case something else weird is going on.

277 \newcommand\sc@defbib [2] {%
278 \renewenvironment*{thebibliography} [1] {%
279  #1*{#2\@mkboth{#2}{#2}}% FIX didn't have this comment char here, OK?
280  \relax\begin{list}{}{%
281  \leftmargin\z@
282  \advance\leftmargin\labelsep
283  \advance\leftmargin\bibindent
284  \itemindent -\bibindent
285  \listparindent \itemindent
286  \parsep \z@}%

The Chicago Manual (University of Chicago Press 1993) does not acknowledge different spacings after different marks of punctuation, distinguish interword from intersentence space, or give rules about where to break a line near an ellipsis. So we are on our own in the bibliography. I have chosen to leave things as they are done in the standard bibliography styles, because I haven’t yet given it my close consideration. That is, we leave all the punctuation the same except for the period, which we set to 1000, I forget now whether that’s a lower or upper case letter. Extending the space after a period when appropriate seems to be the purpose of using \newblock, in this bibstyle.

287  \sfcode`\.=\@m
288  \def\newblock {%
289  \hskip .11em \@plus.33em \@minus.07em%

290 }%

291  \let\@biblabel\Gobble
292  \sloppy
293  \clubpenalty4000\widowpenalty4000%
294  }{%
295  \def\@noitemerr {%
296  \@latex@warning{Empty `thebibliography' environment}%

297 }%

298  \relax\end{list}%

299 }% 300 }

301 \@ifclassloaded{article} {%
302  \sc@defbib{\section}{\refname}%
303 }{% ELSE
304 \sc@defbib{\chapter}{\bibname}%

305 }

306 \InitCS\sc@defbib  % FIX -- where else can I do this?

References

University of Chicago Press. 1993. The Chicago Manual of Style. 14th ed. Chicago: University of Chicago Press.

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\. . . . . . . . . . . . . . . 287 \@auxout . . . . . . . . . 255

\@biblabel . . . . . . . 291

\@cite . . . . . . . 125, 135, 145, 154, 165, 173, 179, 187, 196, 204, 211, 218, 228, 261

\@for . . . . . . . . . . . . 263

\@ifclassloaded . . . 301
\@ifnextchar . . . . . . 246
\@ifundefined . . 11, 267
\@latex@warning . . . 296

\@m . . . . . . . . . . . . . 287 \@minus . . . . . . . . . . 289 \@mkboth . . . . . . . . . 279 \@nameuse . . . . . . . . 271

\@noitemerr . . . . . . . 295
\@openbib@code . . . . 48

\@plus . . . . . . . . . . . 289

\@sc@duplicatedash@false

. . . . . . . . . . . . 66

\@sc@duplicatedash@true

. . . . . . . . . 60, 80

\@tempswafalse . . . . 250
\@tempswatrue . . . . . 247
\@unexpandable@protect

. . . . . . . . . . . 162 \@warning . . . . . . . . 269

\@writeaux . . . 254, 259

\␣ . . . . . . . . 40, 129, 139, 149, 158, 182, 191, 199, 266

A

\advance . . 49, 282, 283

\afterpage . . . . . . . 64

\annotate . . . . . . . . 26

\Annotatefalse . . . . 30
\Annotatetrue . . . . . 27
\AtEndDocument . . . . 55

B \begin . . 40, 95, 103, 280 \bfseries . . . . . . . . 268

                          \bibindent . . . . . . .

49, 50, \underline{276}, 283, 284 \bibitem 61, 67, 69, 70, 72

                          \bibliographystyle . 56

\bibname . . . . . . . . . 304

C \chapter . . . . . . . . . 304 \citation . . . . . . . . 259 \cite . . . . . . . . . . . . \underline{124} \citeA . . . . . . . . . . . \underline{124} \citeANP . . . . . . . . . \underline{124} \citeN . . . . . . . . . . . \underline{124} \citeNP . . . . . . . . . . \underline{124} \citework . . . . . . . . 42 \citeyear . . . . . . . . \underline{216}

                          \citeyearNP . . . . . . . \underline{216}
                          \clubpenalty . . . . . . 293

\csname . . . . . . . . 15, 34

D \def . . . . . . . . . . . 1–

6, 124, 266, 288, 295

                          \defcommand . . . . 48,

125, 128, 135, 138, 145, 148, 154, 157, 165, 173, 179, 181, 187, 190, 196, 198, 204, 211, 218, 221, 228, 231

\do . . . . . . . . . . . . . 263 \docdate . . . . . . . . . . \underline{1}

                          \DoXPackageS . . . . . . . 2
                          \DoXUsepackagE . . . . . \underline{1}

E \edef . . . . . . . . . . . . . 7 \eExpand . . . . . . . . . 15 \else . . 77, 104, 110, 197 \end …. 37, 98, 109, 298

                                                   \endcsname . . . . . 15, 40

\endinput . . . . . . . . 20 environments:

                                                       SCannotation . . . . \underline{8}
                                                       thebibliography    \underline{13}

F \fi 79, 106, 112, 126,

136, 146, 155, 166, 174, 180, 188, 197, 205, 212, 219, 229, 260

\filedate . . . . . . . . . \underline{1} \fileinfo . . . . . . . . . \underline{1}

                                                   \fileversion . . . . . . . \underline{1}

\foo . . . . . . . . . . . 63, 65

G \global . . . . . . . . 66, 67 \Gobble . . . . . . . . . . 291

H

                                                   \HaveECitationS . . . . \underline{1}

\hskip . . . . . . . . . . . 289

I

                                                   \if@filesw . . . . . . . 258
                                                   \if@sc@duplicatedash@

. . . . . . . . . . . . 75

                                                   \if@tempswa . . . . . . .

. 126, 136, 146, 155, 166, 174, 180, 188, 197, 205, 212, 219, 229

                                                   \ifAnnotate . . . 101, 108
                                                   \immediate . . . . . . . 255

\In . . . . . . . . . . . . . 15 \InitCS . . . . . . 262, 306

                                                   \initelyHavECitationS

. . . . . . . . . . . . . 3

                                                   \itemindent . . . . . . .

. . 50, 51, 284, 285

J

\JusTLoaDInformatioN 13

L

\labelsep . . . . . . . . 282

\leftmargin . 49, 281–283

\let . . 54, 67, 69, 72, 85, 89, 102, 168, 176, 207, 214, 238, 242, 265, 291

\listparindent . 51, 285

M

\makeatletter . . . . . 10
\makeatother . . . . . . 19
\MDSavedbibitem . . . 69
\MidCiteWork . . . . 39, 43

N

\NeedsTeXFormat . . . 22

\newblock . . . . . 54, 288

\newboolean . . . . . 25, 59
\newcommand . . 26, 29,

33, 36, 39, 42, 45, 62, 74, 83, 88, 91, 94, 97, 116–118, 121, 134, 144, 153, 164, 172, 178, 186, 195, 203, 210, 217, 227, 237, 241, 245, 277

\newenvironment . . . 100

\newlet . . . . . . . 87, 162 \NewName . . . . . . . . . 257

\noannotate . . . . . . . 29

P

\parsep . . . . . . . 52, 286 \penalty . . . . . . . . . 40

\PostCite . . . . \underline{118},

126, 149, 155, 163, 166, 180, 197, 205, 216, 219

\PostCiteWork . . . \underline{32}, 43

\PPOptArg . . . . . . . \underline{1}, 23

\PreAnnotation . . 45, \underline{94}

\PreChunk . . . . . . . . 46 \PreCite . . . . . \underline{118},

126, 149, 158, 163, 166, 180, 199, 205, 216, 219

\PreCiteWork . . . . \underline{32}, 43
\PreQuotation . . . . . 102
\providecommand . . . 254
\providesavebox . . . 32
\ProvidesPackage . . 23

R

\refname . . . . . . . . . 302 \relax … 54, 95, 98,

103, 109, 280, 298

\renewenvironment . 278
\RequirePackage . 24, 58
\ReserveCS . . . 114, 236
\RestoreDoXVarS . . . 18

\rule . . . . . . . . . . . . 92

S

\SaveCS . . . . . . . . . . 61

\SaveDoXVarS . . . . . . 14

\sbox . . . . . . . . . . . . 98 \sc@@cite …. 248, 251

\sc@begingobble . . . \underline{94}
\sc@bibitemfirst . .

. . . . . . . 62, 67, 72

\sc@box@a . . . . . . . .

. 32, 37, 40, 95, 98 \sc@cite . . 239, 243, \underline{245}

\sc@cite@comma . . . .

…. 151, 160, 201, 224, 234, \underline{236}

\sc@cite@sc 131, 141,

169, 176, 184, 193, 208, 214, \underline{236}

\sc@citeformat . . . .

…. 133, 143, 163, 171, 216, 226

\sc@citesep \underline{236}, 265, 266
\sc@defbib . . . . . . .

. 277, 302, 304, 306

\sc@duplicate@arg 83, 89
\sc@duplicate@dash .

. . . . . . . 74, 85, 87

\sc@endgobble . . . . . \underline{94}
\sc@label . . . . . . . .

. 263, 267, 269, 271

                          \sc@mmmdash . . . . . 76, \underline{91}
                          \sc@one@three . . . . .

…. \underline{114}, 168, 176 \sc@t@a … 262, 264–266

                          \sc@two@three . . . . .

…. \underline{114}, 207, 214

                          SCannotation (environ-

ment) . . . . . . . \underline{94} \SCcite . . . . . . \underline{114},

128, 138, 148, 157, 168, 176, 181, 190, 198, 207, 214, 221, 231

                          \SCduplicate . . . . . . \underline{58}

\section . . . . . . . . . 302

                          \setlength . . . . . . . 276

\sfcode . . . . . . . . . . 287

                          \shortcite . . . . . . . \underline{178}
                          \shortciteA . . . . . . . \underline{178}
                          \shortciteANP . . . . . \underline{178}
                          \shortciteN . . . . . . . \underline{178}
                          \shortciteNP . . . . . . \underline{178}

\sloppy . . . . . . . . . . 292 \space . . . . . . . . . 8, 269 \string . . . . . . . . . . 259

T \the . . . . . . . . . . . 63, 65

                          thebibliography (envi-

ronment) … \underline{277} \thepage . . . . . . . . . 269 \topitem . . . . . . . . . 88 \typeout . . . . . . . 63, 65

U

                          \UndefineCS . . . . . . . 13
                          \UnexpandableProtect

. . . . . . . . . . . 162 \usebox . . . . . . . . . . 37

                          \usepackage . . . . . . . 16

W

                          \widowpenalty . . . . . 293

\write . . . . . . . . . . . 255

Z \z@ . . . . . . 52, 281, 286